# On generalized Turán problems for expansions

### v1

@@ 1 Introduction
Given two graphs $H$ and $F$, let $\mathrm{ex}(n,H,F)$ denote the maximum number of copies of $H$ in an $n$-vertex $F$-free graph. Alon and Shikhelman [1] initiated the systematic study of the function $\mathrm{ex}(n,H,F)$, which is often called the generalized Turán problem. In the case $H=K_{s}$, Zykov [53] and independently Erdős [14] determined the exact value of $\mathrm{ex}(n,K_{s},K_{\ell})$. Wang [49] determined the exact value of $\mathrm{ex}(n,K_{s},M_{k+1})$. In [9], Chakraborti and Chen determined the exact value of $\mathrm{ex}(n,K_{s},P_{\ell})$. Given $r$-graphs $\mathcal{H}$} and $\mathcal{G}$}, let $\mathcal{N}({\mathcal{H}},{\mathcal{G}})$ denote the number of copies of $\mathcal{H}$} in $\mathcal{G}$}.
@@
To the best of our knowledge, besides the results mentioned above, the only other generalized Turán result on hypergraphs is that concerning complete multipartite $r$-graphs, due to Basu, Rödl and Zhao [3]. Specifically, they showed that for $r\geq 2$ and $a_{1}\leq\cdots\leq a_{r+1}$,
$\mathrm{ex}_{r}\big(n,K_{r+1}^{(r)},K_{r+1}^{(r)}(a_{1},\dots,a_{r+1})\big)=o\big(n^{r+1-1/(a_{1}\cdots a_{r})}\big),$
where $K_{r+1}^{(r)}(a_{1},\dots,a_{r+1})$ denotes the complete $(r+1)$-partite $r$-graph with $a_{1},\dots,a_{r+1}$ vertices in its parts.
@@ 2.2 The degenerate case
Thus, ${\mathcal{H}}^{r-1}$ is $P_{3\ell^{\prime}+2}^{r-1}$-free. By Theorems 1.7 and 1.10, we have ${\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})=O(n^{r-2})$. By the definitions of ${\mathcal{S}}^{r-1}$ and ${\mathcal{H}}^{r-1}$, it follows that $|X^{\prime\prime}_{j}|\leq{\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})$, since each $K_{j}^{(r)}\in X^{\prime\prime}_{j}$ corresponds to a copy of $K_{j}^{(r-1)}$ in ${\mathcal{H}}^{r-1}$. Thus, $|X^{\prime\prime}_{j}|=O(n^{r-2})$, which contradicts $|X^{\prime\prime}_{j}|=\Omega(n^{r-1})$. This completes the proof of Claim 5.2. ∎
@@ 6.2 A general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$
In Section 3, we establish a general upper bound on ${\rm{ex}}_{r}(n,K_{s}^{(r)},F^{r})$ for an arbitrary graph $F$ (see Lemma 3.1). Theorems 1.1 and 2.2 show that this bound is sharp in a specific case. Next, we present a more general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$ for any $r$-graph $\mathcal{F}$}. Given a graph $F$, an $r$-graph $\mathcal{H}$} is a Berge-$F$ if there is a bijection $\phi:E(F)\rightarrow E({\mathcal{H}})$ such that $e\subseteq\phi(e)$ for each $e\in E(F)$. Berge [4] defined the Berge-cycle, and Győri, Katona and Lemons [34] defined the Berge-path. Later, Gerbner and Palmer [31] extended the definition to arbitrary graphs $F$. A brief survey of Turán-type results for Berge-graphs can be found in Subsection 5.2.2 in [33]. Here, we further generalize the definition of Berge-graphs to arbitrary $r$-graphs. For $s\geq r$ and an $r$-graph $\mathcal{F}$}, an $s$-graph $\mathcal{G}$} is a Berge-$\mathcal{F}$} if there is a bijection $\phi:E({\mathcal{F}})\rightarrow E({\mathcal{G}})$ such that $e\subseteq\phi(e)$ for each $e\in E({\mathcal{F}})$. Note that for a fixed $r$-graph $\mathcal{F}$}, there are many $s$-graphs that form a Berge-$\mathcal{F}$}. For convenience, we refer to this collection of $s$-graphs as “Berge-$\mathcal{F}$}”. The maximum number of hyperedges in an $n$-vertex $s$-graph containing no subhypergraph isomorphic to any Berge-$\mathcal{F}$} is denoted by ${\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})$.
The problem of counting copies of the complete $r$-graph $K_{s}^{(r)}$ in an $n$-vertex $\mathcal{F}$}-free $r$-graph is closely related to the study of Berge-hypergraphs. In what follows, we establish an explicit relationship between these two classes of problems.
Proposition 6.3.

### v2

@@ 1 Introduction
Given two graphs $H$ and $F$, let $\mathrm{ex}(n,H,F)$ denote the maximum number of copies of $H$ in an $n$-vertex $F$-free graph. Alon and Shikhelman [1] initiated the systematic study of the function $\mathrm{ex}(n,H,F)$, which is often called the generalized Turán problem. In the case where $H=K_{s}$, Zykov [60] and independently Erdős [17] determined the exact value of $\mathrm{ex}(n,K_{s},K_{\ell})$. Wang [55] determined the exact value of $\mathrm{ex}(n,K_{s},M_{k+1})$. In [12], Chakraborti and Chen determined the exact value of $\mathrm{ex}(n,K_{s},P_{\ell})$. Given $r$-graphs $\mathcal{H}$} and $\mathcal{G}$}, let $\mathcal{N}({\mathcal{H}},{\mathcal{G}})$ denote the number of copies of $\mathcal{H}$} in $\mathcal{G}$}.
@@
Balko, Gerbner, Kang, Kim and Palmer [3] established the asymptotics of the generalized Turán number for expansions of graphs $F$ with $\chi(F)>s>r$.
Theorem 1.10 (Balko, Gerbner, Kang, Kim and Palmer [3]).
Let $F$ be a graph and let $\chi(F)=k>s>r$. Then
$\mathrm{ex}_{r}(n,K_{s}^{(r)},F^{r})=(1+o(1))\cdot e({\mathcal{T}}_{s}(n,k-1)).$
@@
For general hypergraphs, Kirsch and Radcliffe [40] established bounds on the number of $s$-cliques in an $r$-graph $\mathcal{H}$ with bounded $i$-degree $\Delta$ (here, the $i$-degree denotes the number of hyperedges containing a given $i$-vertex set), in three settings: $\mathcal{H}$ has $n$ vertices; $\mathcal{H}$ has $m$ hyperedges; and $\mathcal{H}$ has a fixed number of $t$-cliques for some $t$ satisfying $r\leq t\leq s$. In particular, for special values of $\Delta$, they characterized the extremal $r$-graphs and showed that the corresponding bounds are tight.
Let us now turn our attention to $r$-graphs that are not expansions. Let $S_{2}^{(3)}$ denote the unique $3$-graph on $4$ vertices with $2$ hyperedges. Bodnár et al. [7] showed that determining the Turán number of a $3$-graph $\mathcal{F}$} in the $\ell_{2}$-norm (see [4] for the relevant definition) is equivalent to maximizing $2{\mathcal{N}}(S_{2}^{(3)},{\mathcal{H}})+3e({\mathcal{H}})$ over all $\mathcal{F}$}-free 33-graphs $\mathcal{H}$}. Their result implies that for sufficiently large $n$,
$\mathrm{ex}_{3}(n,S_{2}^{(3)},K_{4}^{(3)})={\mathcal{N}}(S_{2}^{(3)},{\mathcal{C}}_{n}^{(3)}),$
where ${\mathcal{C}}_{n}^{(3)}$ denote the $3$-graph on $[n]$ consisting of all triples of types $V_{1}V_{2}V_{3}$, $V_{1}V_{1}V_{2}$, $V_{2}V_{2}V_{3}$ and $V_{3}V_{3}V_{1}$ for a balanced partition $V_{1},V_{2},V_{3}$ of $[n]$.
For $k\geq r$, let $K_{\ell_{1},\ell_{2},\dots,\ell_{k}}^{(r)}$ denote the complete $k$-partite $r$-graph with $\ell_{1},\ell_{2},\dots,\ell_{k}$ vertices in its parts. Ma, Yuan and Zhang [47] proved that for any $r$-graph $\mathcal{H}$}, there exists a constant $c>0$ such that for any integer $p\geq c$,
$\mathrm{ex}_{r}(n,{\mathcal{H}},K_{s_{1},s_{2},\dots,s_{r-1},p}^{(r)})=\Omega\left(n^{|V({\mathcal{H}})|-\frac{e({\mathcal{H}})}{s_{1}s_{2}\cdots s_{r-1}}}\right).$
In particular, they established that for any positive integers $a_{1},a_{2},\dots,a_{r},s_{1},s_{2},\dots,s_{r-1}$ and $r\geq 2$ satisfying that $a_{1}<s_{1}\leq s_{2}\leq\dots\leq s_{r-1}$ and $a_{i}\leq s_{i-1}$ for each $2\leq i\leq r$, there exists a constant $c>0$ such that for any integer $p\geq c$,
$\mathrm{ex}_{r}(n,K_{a_{1},a_{2},\dots,a_{r}}^{(r)},K_{s_{1},s_{2},\dots,s_{r-1},p}^{(r)})=\Theta\left(n^{\left(a_{1}+a_{2}+\dots+a_{r}\right)-\frac{a_{1}a_{2}\cdots a_{r}}{s_{1}s_{2}\cdots s_{r-1}}}\right).$
Recently, Basu, Rödl and Zhao [5] showed that for any positive integers $a_{1}\leq a_{2}\leq\cdots\leq a_{r+1}$ and $r\geq 2$,
$\mathrm{ex}_{r}\big(n,K_{r+1}^{(r)},K_{a_{1},a_{2},\dots,a_{r+1}}^{(r)}\big)=o\left(n^{r+1-\frac{1}{a_{1}a_{2}\cdots a_{r}}}\right).$
To the best of our knowledge, besides the results mentioned above, the only other generalized Turán result on hypergraphs is that concerning complete bipartite $r$-graphs, which was introduced by Mubayi and Verstraëte [49].
Definition 1.12 (Complete bipartite $r$-graph).
Let $X_{1},X_{2},\dots,X_{t}$ be $t$ pairwise disjoint sets of size r1r-1, and let $Y$ be a set of $s$ elements, disjoint from $\bigcup_{i=1}^{t}X_{i}$. Then $K_{s,t}^{(r)}$ denotes the $r$-graph with vertex set $\bigcup_{i=1}^{t}X_{i}\cup Y$ and hyperedge set $\left\{X_{i}\cup\{y\}:i\in[t],\;y\in Y\right\}$.
Note that $K_{s,t}^{(r)}$ is a hypergraph extension of the complete bipartite graph. In [49], Mubayi and Verstraëte obtained some bounds on $\mathrm{ex}_{r}(n,K_{s,t}^{(r)})$ for $s\leq t$. Subsequently, Xu, Zhang and Ge [56] proved the following result.
Theorem 1.13 (Xu, Zhang and Ge [56]).
Let $r\geq 3$. For any positive integer $t$ and any $r$-graph $\mathcal{H}$, there exists some constant $c$ which depends on $r$ and $t$ such that if $s\geq c$, then
$\mathrm{ex}_{r}(n,\mathcal{H},K_{s,t}^{(r)})=\Omega\left(n^{|V({\mathcal{H}})|-\frac{e({\mathcal{H}})}{t}}\right).$
Furthermore, they showed that there exists some constant $c^{\prime}_{r,t}$ which depends on $r$ and $t$ such that if $s\geq c^{\prime}_{r,t}$ and $b<t$,
$\mathrm{ex}_{r}(n,K_{1,b}^{(r)},K_{s,t}^{(r)})=\Theta\left(n^{1+b(r-1)-\frac{b}{t}}\right).$
@@ 2.2 The degenerate case
Thus, ${\mathcal{H}}^{r-1}$ is $P_{3\ell^{\prime}+2}^{r-1}$-free. By Theorems 1.7 and 1.11, we have ${\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})=O(n^{r-2})$. By the definitions of ${\mathcal{S}}^{r-1}$ and ${\mathcal{H}}^{r-1}$, it follows that $|X^{\prime\prime}_{j}|\leq{\mathcal{N}}(K_{j}^{(r-1)},{\mathcal{H}}^{r-1})$, since each $K_{j}^{(r)}\in X^{\prime\prime}_{j}$ corresponds to a copy of $K_{j}^{(r-1)}$ in ${\mathcal{H}}^{r-1}$. Thus, $|X^{\prime\prime}_{j}|=O(n^{r-2})$, which contradicts $|X^{\prime\prime}_{j}|=\Omega(n^{r-1})$. This completes the proof of Claim 5.2. ∎
@@ 6.2 A general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$
In Section 3, we establish a general upper bound on ${\rm{ex}}_{r}(n,K_{s}^{(r)},F^{r})$ for an arbitrary graph $F$ (see Lemma 3.1). Theorems 1.1 and 2.2 show that this bound is sharp in a specific case. Next, we present a more general bound on $\mathrm{ex}_{r}(n,K_{s}^{(r)},{\mathcal{F}})$ for any $r$-graph $\mathcal{F}$}. Given a graph $F$, an $r$-graph $\mathcal{H}$} is a Berge-$F$ if there is a bijection $\phi:E(F)\rightarrow E({\mathcal{H}})$ such that $e\subseteq\phi(e)$ for each $e\in E(F)$. Berge [6] defined the Berge-cycle, and Győri, Katona and Lemons [37] defined the Berge-path. Later, Gerbner and Palmer [34] extended the definition to arbitrary graphs $F$. A brief survey of Turán-type results for Berge-graphs can be found in Subsection 5.2.2 in [36]. Subsequently, Balko et al. [3] further generalized the definition of Berge-graphs to arbitrary hypergraphs. For $s\geq r$ and an $r$-graph $\mathcal{F}$}, an $s$-graph $\mathcal{G}$} is a Berge-$\mathcal{F}$} if there is a bijection $\phi:E({\mathcal{F}})\rightarrow E({\mathcal{G}})$ such that $e\subseteq\phi(e)$ for each $e\in E({\mathcal{F}})$. Note that for a fixed $r$-graph $\mathcal{F}$}, there are many $s$-graphs that form a Berge-$\mathcal{F}$}. For convenience, we refer to this collection of $s$-graphs as “Berge-$\mathcal{F}$}”. The maximum number of hyperedges in an $n$-vertex $s$-graph containing no subhypergraph isomorphic to any Berge-$\mathcal{F}$} is denoted by ${\rm{ex}}_{s}(n,{\rm Berge}{\text{-}}{\mathcal{F}})$.
The problem of counting copies of the complete $r$-graph $K_{s}^{(r)}$ in an $n$-vertex $\mathcal{F}$}-free $r$-graph is closely related to the study of Berge-hypergraphs. An explicit relationship between these two classes of problems was first established by Balko et al. [3], and we provide an alternative proof here.
Proposition 6.3 (Balko, Gerbner, Kang, Kim and Palmer [3]).
@@
Acknowledgements
We would like to express our sincere gratitude to Professor Dániel Gerbner for his helpful comments on an earlier draft of this paper.